# Learning Molecular Representation in a Cell

### v2

@@ 5.3.1 Experimental Setting
For InfoAlign, we use the pretrained decoder from Section 5.1 to extract the morphology features of the encoded molecule and then calculate the likelihood of these decoded features against the candidate morphology data. We then rank the candidates in the decoding space based on their likelihood scores.
@@
5.3.2 Results and Analysis
5.4 RQ3: Hyperparameter Analysis
@@ Appendix A Proof Details
For the molecule $X$, its latent representation $Z$, and any biological modality variable $Y$, the exact computation of $I(Z;Y)$ and $I(X;Z)$ is intractable due to the unknown conditional distribution $p(Y|Z)$ and the marginal $p(Z)$. We introduce the variational approximations $q(Y|Z)$ for a lower bound on $I(Z;Y)$:
@@
For the molecular representation $Z$, and any biological modality variable $Y$, we use the neural network parameterized critic $h(Z,Y)$ with the energy-based variational family for density approximation [37]:
@@
Suppose there are $K-1$ additional examples independently and identically sampled from the modality of $X$, denoted as $Y_{2:K}$, and the critic is configured with parameters $a(\cdot)$ as $1+\log\frac{e^{h(Z,Y)}}{a(Z;Y,Y_{2:K})}$. Then, we can rewrite $I_{NWJ}$ for its multi-sample version:
@@ Appendix B Context Graph Details
Figure 5: From the initial idea in Section 4 to the practical implementation of the context graph, we first display relations between molecules and all the landmark genes from [54] for the $X_{1}-E_{3}$ and $X_{3}-E_{2}$ relationships. $E_{3}$ and $E_{2}$ are landmark genes involved in small molecule perturbations and cell morphology perturbation; we display them separately for clarity. Next, we merge all landmark genes into new gene expression nodes and integrate genes from genetic perturbations in the JUMP dataset [6] with cell morphology features. Practical considerations are detailed in Sections 5.1 and B.
@@ B.1 Edge Construction
    $e\in\mathcal{Y}^{[X]}\times\mathcal{Y}^{[C]}$: These edges are introduced through molecule perturbation experiments from cell painting datasets created by [3] and the JUMP dataset [6]. It links molecule nodes with cell morphology nodes. We use the edge weight $1$ for all these edges.
@@
    $e\in\mathcal{Y}^{[E]}\times\mathcal{Y}^{[C]}$: These edges should be introduced by genetic perturbation from the JUMP dataset [6]. The perturbations are either based on gene overexpression (ORF) or gene knockout techniques (CRISPR). They link the gene nodes and the cell morphology nodes. However, the genes introduced by the genetic perturbations lack gene expression profiling from [49] as node features. We did not implement gene-cell morphology edges from [6] due to the absence of differential gene expression profiling values [49]. Instead, we merged the gene nodes from [6] with their linked cell morphology nodes, creating single nodes. This approach enables a more efficient context graph, incorporating some gene nodes with cell morphology features.
@@
    $e\in\mathcal{Y}^{[X]}\times\mathcal{Y}^{[E]}$: These edges could represent molecule-gene binding and regulation relationships, linking molecules to genes [16]. Some links can be sourced from [16], and we also retrieve gene-molecule links from [54] by selecting the top 5% absolute differential expression values.
@@
    $e\in\mathcal{Y}^{[E]}\times\mathcal{Y}^{[E]}$: These edges denote the relationships of gene-gene covariance and interaction and we use the links from [16].
@@ Appendix C Experiment Details
Table 5: Complete results on ChEMBL2K. We report average AUC (Avg.), as well as the percentage of tasks achieving AUC above 80%, 85%, 90% and 95%. The best results in each category are highlighted and reported in Table 2.
Repr. Model Avg. > 80 % > 85 % > 90 % > 95% Morgan FP MLP 76.8±2.2 48.8±3.9 34.6±6.3 21.9±5.7 12.2±2.4 RF 54.7±0.7 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 GP 51.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 AttrMask 74.0±0.5 46.8±2.7 31.2±4.4 14.6±1.7 2.4±0.0 Pretrained GNN ContextPred 77.0±0.5 55.1±1.3 34.2±4.6 14.6±1.7 1.5±1.3 EdgePred 75.6±0.5 54.1±4.0 34.6±7.2 12.2±2.4 6.3±1.3 Pretrained ChemLM Roberta-102M 74.7±1.9 46.3±3.5 35.1±4.4 22.9±1.3 7.8±3.2 GPT2-87M 71.0±3.4 31.2±11.2 20.0±9.4 7.3±6.9 2.9±3.2 MolT5 70.0±0.8 32.2±2.0 21.0±4.1 8.8±1.3 3.4±1.3 ChemGPT-19M 65.0±1.1 16.1±2.8 11.2±3.3 5.4±1.1 2.9±1.1 Cell Morphology MLP 64.3±2.4 15.6±6.6 8.3±3.7 4.9±3.9 2.9±2.7 RF 55.9±0.7 3.9±1.3 3.9±1.3 2.4±0.0 2.4±0.0 GP 51.5±0.0 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 Gene Expression MLP 56.1±1.1 5.1±1.4 3.4±1.3 3.4±1.3 2.4±0.0 RF 52.8±0.3 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 GP Run out of time Aligned CLOOME 66.7±1.8 26.8±4.6 16.1±3.7 10.7±5.1 3.4±2.8 InfoCORE (GE) 79.3±0.9 62.4±2.8 46.3±3.0 30.3±2.2 18.1±1.3 InfoCORE (CP) 73.8±2.0 37.6±9.2 26.3±4.7 10.7±4.1 5.9±2.2 InfoAlign 81.3±0.6 66.3±2.7 49.3±2.7 35.1±3.7 18.5±2.8
Figure 6: An overview of representation’s predictive performance on all 41 bioactivity prediction tasks on ChEMBL2K.
Table 6: Complete results on Broad6K. We report average AUC (Avg.), as well as the percentage of tasks achieving AUC above 80%, 85%, 90% and 95%. The best results in each category are highlighted and reported in Table 2.
Repr. Model Avg. > 80 % > 85 % > 90 % > 95% Morgan FP MLP 63.3±0.3 6.3±0.0 4.4±1.7 3.1±0.0 3.1±0.0 RF 55.5±0.1 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 GP 50.6±0.0 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 AttrMask 59.8±0.2 3.1±0.0 3.1±0.0 3.1±0.0 0.0±0.0 Pretrained GNN ContextPred 60.0±0.2 7.5±1.7 3.1±0.0 3.1±0.0 0.0±0.0 EdgePred 59.9±0.2 3.1±0.0 3.1±0.0 3.1±0.0 0.0±0.0 Roberta-102M 59.8±0.7 5.0±1.7 3.1±0.0 3.1±0.0 0.0±0.0 Pretrained ChemLM GPT2-87M 60.6±0.3 7.5±1.7 3.1±0.0 1.9±1.7 0.0±0.0 MolT5 56.4±0.8 3.8±1.4 2.5±1.4 2.5±1.4 0.0±0.0 ChemGPT-19M 55.1±0.9 3.1±0.0 3.1±0.0 1.3±1.7 0.0±0.0 MLP 51.9±1.0 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 Cell Morphology RF 55.3±0.1 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 GP 54.7±0.0 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 Gene Expression MLP 56.9±1.4 1.9±1.7 1.9±1.7 1.9±1.7 0.0±0.0 RF 55.2±0.2 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 GP 50.1±0.0 0.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 Aligned CLOOME 61.7±0.5 3.1±0.0 3.1±0.0 0.0±0.0 0.0±0.0 InfoCORE (GE) 60.2±0.2 3.1±0.0 0.0±0.0 0.0±0.0 0.0±0.0 InfoCORE (CP) 61.1±0.2 6.3±0.0 3.1±0.0 0.0±0.0 0.0±0.0 InfoAlign 69.9±0.1 18.8±2.2 3.1±0.0 3.1±0.0 1.3±1.7
Figure 7: An overview of representation’s predictive performance on five major task categories on Broad6K.
Table 7: Complete results on ToxCast. We report average AUC (Avg.), as well as the percentage of tasks achieving AUC above 80%, 85%, 90% and 95%. The best results in each category are highlighted and reported in Table 3.
Representation Models Avg. >80 % >85 % >90 % >95 % Morgan FP MLP 57.6±1.0 1.6±0.3 0.6±0.4 0.3±0.3 0.0±0.0 RF 52.3±0.1 0.2±0.1 0.1±0.1 0.0±0.1 0.0±0.0 GP Run out of time AttrMask 63.1±0.8 3.2±1.2 0.8±0.3 0.2±0.1 0.0±0.1 Pretrained GNN ContextPred 63.0±0.6 3.3±1.3 0.4±0.4 0.0±0.0 0.0±0.0 EdgePred 63.5±1.1 4.8±3.0 0.8±0.4 0.1±0.1 0.0±0.0 Roberta-102M 64.2±0.8 3.1±1.8 0.9±0.6 0.2±0.2 0.0±0.1 Pretrained ChemLM GPT2-87M 61.5±1.1 2.4±0.6 0.8±0.3 0.3±0.1 0.0±0.1 MolT5 64.7±0.9 3.6±1.1 1.1±0.3 0.4±0.2 0.0±0.1 ChemGPT-19M Token Error Aligned CLOOME 54.2±1.0 0.9±0.2 0.3±0.1 0.1±0.1 0.0±0.0 InfoCORE (GE) 65.3±0.2 5.4±1.7 1.2±0.4 0.3±0.1 0.0±0.0 InfoCORE (CP) 62.4±0.4 1.3±0.5 0.3±0.3 0.0±0.0 0.0±0.0 InfoAlign 66.4±1.1 6.6±1.6 2.1±0.5 0.7±0.3 0.1±0.1
Table 8: Complete results on Biogen3K. We report the average MAE and MAE for all tasks (scaled by $\times$ 100). The best results in each category are highlighted and reported in Table 3.
Representation Models Avg. hPPB rPPB RLM HLM ER Solubility MLP 66.2±2.4 66.1±2.6 56.8±2.3 56.5±4.2 74.6±6.2 73.7±7.3 69.5±3.0 Morgan FP RF 52.8±0.2 44.2±0.1 44.2±0.1 41.9±0.2 67.7±0.7 66.9±0.9 51.6±0.1 GP 60.0±0.0 51.3±0.0 59.5±0.0 49.7±0.0 68.8±0.0 69.3±0.0 61.6±0.0 Pretrained GNN AttrMask 67.3±0.3 82.4±1.1 49.8±0.7 51.7±1.0 57.9±0.6 62.6±0.5 99.1±1.2 ContextPred 68.5±0.9 85.0±7.9 49.7±0.4 55.1±2.7 61.4±1.8 63.1±0.5 96.5±3.7 EdgePred 67.8±0.9 81.2±10.2 48.0±0.5 53.5±2.8 62.2±1.8 62.9±0.7 99.1±6.9 Roberta-102M 69.0±2.6 71.4±14.5 65.1±19.2 63.7±24.6 67.5±5.2 69.9±4.9 76.7±13.2 Pretrained ChemLM GPT2-87M 73.9±8.5 65.4±12.9 73.1±20.8 54.1±12.9 83.2±21.5 86.1±19.8 81.8±25.5 MolT5 65.1±0.5 76.7±2.1 55.9±1.1 49.2±1.0 70.3±0.8 73.1±1.0 65.3±1.7 ChemGPT-19M 75.7±8.5 59.5±7.3 88.8±32.3 76.1±11.8 84.0±20.6 77.2±8.5 68.6±7.1 Aligned CLOOME 64.3±0.4 65.2±1.5 56.9±0.8 44.2±0.8 70.7±0.4 73.6±0.8 75.0±2.1 InfoCORE (GE) 69.9±1.2 79.9±3.6 51.6±1.8 51.3±2.1 78.6±0.3 77.8±1.9 80.4±0.9 InfoCORE (CP) 71.0±0.6 74.5±4.9 53.5±0.7 53.6±2.1 80.8±1.5 79.4±3.4 84.4±1.0 InfoAlign 49.4±0.2 39.7±0.4 39.2±0.3 40.5±0.6 66.7±1.7 62.0±1.5 48.4±0.6
@@ C.2 Implementation and Baseline
    Pretrained GNN representations [17] (Structure only): We consider AttrMask, ContextPred, and EdgePred with supervised pretraining [17] and apply fine-tuned MLPs on top of the pretrained representations.
@@
    Pretrained ChemLM representations [12] (Structure only): We consider pretrained models such as 102M Roberta, 87M GPT2, MolT5, and 19M ChemGPT [12]. We apply fine-tuned MLPs on top of these pretrained representations.
@@ C.3 More Results for Molecular Property Prediction
We present comprehensive results for the ChEMBL2K dataset, including 19 baseline performances in Table 5 and comparisons of basic representation approaches with the InfoAlign across all task dimensions in Figure 6. Similarly, results for the Broad6K dataset, comparing InfoAlign against 19 baselines are in Table 6 and comparison with basic representations across five major task dimensions (Cell, Yeast, Viral, Biochem, and Bacterial related targets) are in Figure 7. Results for the ToxCast and Biogen3K datasets are provided in Table 7 and Table 8, respectively. From these detailed results we have more observations:
@@
(2) Learning universal molecular representations solely from molecular structures remains challenging, even within the representation category. For pretrained GNN representations, ContextPred outperforms others in the ChEMBL2K and Broad6K datasets but not in other datasets. For ChemLM representations, MolT5 excels over other sequential-based models in the ToxCast and Biogen3K datasets, but this is not the case with the ChEMBL2K and Broad6K datasets. Different datasets may emphasize varied aspects of bioactivity classification or regression and pose generalization challenges for molecular representation learning.

### v3

@@ 6.2.1 Experimental Setting
For InfoAlign, we use the pretrained decoder from Section 5 to extract the morphology features of the encoded molecule and then calculate the likelihood of these decoded features against the candidate morphology data. We then rank the candidates in the decoding space based on their likelihood scores.
6.2.2 Results and Analysis
6.3 RQ3: Performance Analysis
6.3.1 Ablation Studies
Table 5: Ablation studies on the pretraining loss. Different node features are removed from the context graph to assess their impact on downstream tasks. Avg. AUC is reported.
ChEMBL2K AUC $\uparrow$ Broad6K AUC $\uparrow$ ToxCast AUC $\uparrow$ Biogen3K MAE $\downarrow$ ($\times$100) Default as Eq. 3 $81.3\pm 0.6$ $70.0\pm 0.1$ $66.4\pm 1.1$ $49.4\pm 0.2$ w/o Cell Morphology $80.7\pm 0.6$ $68.6\pm 0.1$ $65.5\pm 1.1$ $51.7\pm 1.1$ w/o Gene Expressions $78.3\pm 0.5$ $68.6\pm 0.2$ $64.7\pm 1.0$ $50.3\pm 0.5$ w/o Molecular Features $79.1\pm 0.2$ $67.1\pm 0.4$ $65.8\pm 2.3$ $51.7\pm 0.6$
We perform ablation studies on Eq. 3 by pretraining encoders with different targets removed: (1) molecule-related, (2) cell morphology-related, and (3) gene expression-related features. The results in Table 5 cover all datasets. We observe that both cell morphology and gene expression features are crucial for achieving the best performance. Different biological targets have varying impacts across datasets: molecular structure has more influence on Broad6K and Biogen3K, while gene expression is more important for ChEMBL2K and ToxCast.
@@ 6.3.1 Ablation Studies
6.3.2 Hyperparameter Analysis
@@ Appendix A Proof Details
For the input, latent, and target variables $X$, $Z$, and $Y$, the exact computation of the mutual information (MI) $I(Z;Y)$ and $I(X;Z)$ is intractable. For the molecule $x$, its latent representation $z$, and any biological target from cellular responses $y$, we introduce variational approximations $q(y|z)$ to obtain a lower bound on $I(Z;Y)$:
@@
For the molecule $x$, its latent representation $z$, and any biological target from cellular responses $y$, we use the neural network parameterized critic $h(z,y)$ with the energy-based variational family for density approximation [38]:
@@
Suppose there are $K-1$ additional examples independently and identically sampled and denoted as $y_{2:K}$, and the critic is configured with parameters $a(\cdot)$ as $1+\log\frac{e^{h(z,y)}}{a(z;y,y_{2:K})}$. Then, we can rewrite $I_{NWJ}$ for its multi-sample version:
@@ Appendix B Context Graph Details
Figure 5: From the initial idea in Section 4 to the practical implementation of the context graph, we first display relations between molecules and all the landmark genes from [56] for the $X_{1}-E_{3}$ and $X_{3}-E_{2}$ relationships. $E_{3}$ and $E_{2}$ are landmark genes involved in small molecule perturbations and cell morphology perturbation; we display them separately for clarity. Next, we merge all landmark genes into new gene expression nodes and integrate genes from genetic perturbations in the JUMP dataset [6] with cell morphology features. Practical considerations are detailed in Sections 5 and B.
@@ B.1 Edge Construction
    Molecule-Cell Morphology Edges: These edges are introduced through molecule perturbation experiments from cell painting datasets created by [3] and the JUMP dataset [6]. It links molecule nodes with cell morphology nodes. We use the edge weight $1$ for all these edges.
@@
    Edge-Cell Morphology Edges: These edges should be introduced by genetic perturbation from the JUMP dataset [6]. The perturbations are either based on gene overexpression (ORF) or gene knockout techniques (CRISPR). They link the gene nodes and the cell morphology nodes. However, the genes introduced by the genetic perturbations lack gene expression profiling from [50] as node features. We did not implement gene-cell morphology edges from [6] due to the absence of differential gene expression profiling values [50]. Instead, we merged the gene nodes from [6] with their linked cell morphology nodes, creating single nodes. This approach enables a more efficient context graph, incorporating some gene nodes with cell morphology features.
@@
    Molecule-Gene Edges: These edges could represent molecule-gene binding and regulation relationships, linking molecules to genes [16]. Some links can be sourced from [16], and we also retrieve gene-molecule links from [56] by selecting the top 5% absolute differential expression values.
@@
    Gene-Gene Edges: These edges denote the relationships of gene-gene covariance and interaction and we use the links from [16].
@@ Appendix C Experiment Details
Figure 6: An overview of the representation’s predictive performance on all 41 bioactivity prediction tasks in ChEMBL2K. Results for molecular structure are obtained from the best method ContextPred. Results for cell morphology and gene expression come from the best method based on MLPs.
Figure 7: An overview of the representation’s predictive performance across five major task categories on Broad6K. Results for molecular structure are obtained from the best method based on JOAO. Results for cell morphology come from the best method based on RF. Results for gene expression are derived from the best method based on MLP.
@@ C.2 Implementation and Baseline
    Pretrained GNN representations [18] (Structure only): We consider AttrMask, ContextPred, and EdgePred with supervised pretraining [18]. We also include GraphCL [60], GROVER [44], JOAO [61], MGSSL [62], GraphLoG [59], GraphMAE [17], DSLA [23], and UniMol [63]. We implement GraphCL, GROVER, and JOAO based on [55]. Fine-tuned MLPs are applied on top of the pretrained representations.
@@
    Pretrained ChemLM representations [12] (Structure only): We consider pretrained models such as 102M Roberta and 87M GPT2 implemented by [31]. We also include MolT5 [10] and 19M ChemGPT [12]. We apply fine-tuned MLPs on top of these pretrained representations.
@@ C.3 More Results for Molecular Property Prediction
We present additional comparisons on the ChEMBL2K dataset between basic representation approaches and InfoAlign across all task dimensions in Figure 6. Similarly, results for the Broad6K dataset, comparing basic representations across five major task dimensions (Cell, Yeast, Viral, Biochem, and Bacterial-related targets), are shown in Figure 7. Combined with Tables 2 and 3, these detailed results lead to further observations:
@@
(2) Learning universal molecular representations solely from molecular structures remains challenging, even within the representation category. For pretrained GNN representations, ContextPred outperforms others on the ChEMBL2K dataset. JOAO excels on the Broad6K dataset. UniMol and GraphMAE are the best pretrained GNN representations on ToxCast and Biogen3K datasets, respectively. For ChemLM representations, MolT5 excels over other sequential-based models in the ToxCast and Biogen3K datasets, but this is not the case with the ChEMBL2K and Broad6K datasets. Different datasets may emphasize varied aspects of bioactivity classification or regression and pose generalization challenges for molecular representation learning.